OpenStudy (anonymous):

How many grams of NaI would be used to produce a 23.0 M solution with a volume of 1.50 L?

OpenStudy (anonymous):

First question: [(3.5mol CaCl2) / 1.0 L] * 2.0 L = 7.0mol CaCl2 7.0mol CaCl2 * (110.98g CaCl2 / 1molCaCl2) = 780g CaCl2 Second question: (Does this mean 5.00 x 102ml = 510mL solution?) We need to convert the solution to liters because molarity is in units of mol/L: 510mL * (1L / 1000mL) = 0.540L Convert the mass of KI to moles: 249g KI * (1mol KI / 166.0g KI) = 1.50mol KI 1.50mol KI / 0.540L = 2.78 M Third question: 1.5L * (2.5mol / 1.0L) = 3.75mol LiF; keeping with sig. figs.: 3.8mol LiF Fourth question: 2.0M solution of NaI = (2.0molNaI / 1 L) So since it already has 1L volume, we need only to convert the moles to grams: 2.0mol NaI * (149.89g NaI / 1mol NaI) = 0.013g NaI Fifth question: (Does this mean 2.50 x 102cm3 = 255cm3 solution?) Note that 1cm3 = 1 mL so we convert the cm3 to Liters: 255cm3 * (255mL / 255cm3) * (1L / 1000mL) = 0.255L of solution. Convert the mass of C6H12o6 to moles: 45.0g C6H12O6 * (1mol C6H12O6 / 180.16g C6H12O6) = 0.250mol C6H12O6. Now: 0.250molC6H12O6 / 0.255L = 0.980M solution. Sixth question: 2.50L * [ (3.5mol Sr(NO3)2) / 1L] = 8.75mol but with sig. figs. it is 8.8mol Sr(NO3)2. Hope this helps!